# Towards deployment-centric multimodal AI beyond vision and language

### v1

@@ Acknowledgements
X.L., J.Z., S.Z. and H.L. contributed equally. X.L., J.Z., S.Z., T.L.vdP., S.T., T.Z., W.K.C., P.H.C., and H.L. conceptualised the manuscript. X.L., J.Z., S.Z., A.G., M.Z., L.vZ., O.T., and H.L. designed the figures and table. X.L. and H.L. coordinated the entire project. All authors contributed to writing, resources, or editing.
@@
Beyond its critical role in pandemic response, multimodal AI holds transformative potential in healthcare by advancing diagnostics, personalising treatment, and improving patient care. By integrating diverse data sources such as electronic health records (EHRs), imaging, biosignals, and multiomics, these systems can provide comprehensive insights into patient health acosta2022multimodal. For example, integrating wearable biosignals with EHRs can enable continuous monitoring, allowing for the early detection of chronic conditions such as arrhythmias and heart failure krittanawong2021integration. In oncology, the fusion of imaging modalities and multiomics data can enhance cancer diagnosis and treatment planning steyaert2023multimodal, offering a more nuanced understanding of disease progression. Similarly, in neurology, combining imaging, sensor data, and clinical observations can facilitate the early detection of neurodegenerative diseases huang2023multimodal.
Wider adoption of multimodal AI is hindered by challenges such as privacy concerns, the heterogeneity of healthcare data, and the scarcity of high-quality datasets. Addressing these challenges requires standardised methodologies, privacy-preserving techniques, and interdisciplinary collaborations. Building datasets that reflect the complexities of patient pathways and addressing bottlenecks in data integration are crucial steps towards integrating multimodal AI into routine clinical practice. Techniques like federated learning, differential privacy, and privacy-preserving generative AI for synthetic multimodal data generation can mitigate privacy risks, enabling the development of robust and trustworthy AI systems. Integrating multimodal data, such as multiomics lunke2023integrated, into routine clinical workflows will require rigorous validation and improved standardisation. Notably, limited validation through randomised controlled trials han2024randomised has led to the emergence of community-based guidelines Chan2020ReportingGF aimed at improving AI reliability and transparency.
@@
In social science, multimodal AI can drive innovations in behavioural analysis, public policy decision-making, and societal impact assessments. Beyond its applications in pandemic response, multimodal AI can help monitor societal behaviours and safeguard vulnerable populations, such as protecting children online, by integrating text, image and metadata from social media ali2023getting. Multimodal AI systems can also help detect online criminal activity goyal2023detection and analyse societal trends to inform public policy androutsopoulou2018framework. Additionally, acoustic data can add depth to communication analysis by capturing emotional nuances zhu2024review, while combining economic transaction data with geospatial information can reveal trends in societal decision-making Huang2024CrowdsourcedGD.
@@
Beyond self-driving cars, multimodal AI supports advanced autonomous systems such as robotics, precision manufacturing, and medical applications. Robots equipped with multimodal AI can integrate vision, haptic feedback, and sensor data to perform delicate tasks, such as in surgical procedures he2022robotic or complex manipulation wang2024multimodal, where precision is paramount. Additionally, multimodal AI can enable seamless human-robot interaction by combining speech recognition, visual data, and environmental context duncan2024survey. Speech technology and natural language processing can enhance user interfaces, improving interaction while minimising distractions for operators in some robotic applications.
Despite the potential of multimodal AI, challenges remain in achieving effective data fusion, system interoperability, and robust model generalisation. Furthermore, the lack of diverse datasets, particularly for edge cases or atypical deployment environments, limits the scalability of these systems. Establishing open platforms for data and toolkit sharing accelerates innovation and ensures interoperability across engineering applications. Integrating insights from legal studies and social sciences can enhance compliance with evolving regulations and improve user-centred design, boosting societal acceptance of autonomous systems. Environmental science can contribute to developing energy-efficient and sustainable designs, particularly for robotics and manufacturing processes. Addressing these challenges requires interdisciplinary collaborations to improve data diversity and develop scalable, interoperable systems.
@@
Multimodal AI is transforming scientific discovery by integrating diverse data sources to model complex phenomena. Beyond its role in vehicle dynamics simulation, as seen in self-driving car design, multimodal AI is accelerating materials science advancements. For instance, integrating experimental and computational data enables the development of improved batteries, fuel cells and supercapacitors liu2024recent. We can incorporate domain knowledge via techniques such as physics-informed neural networks raissi2019physics; karniadakis2021physics and neurosymbolic AI garcez2023neurosymbolic to improve the accuracy of scientific simulations and enhance model trustworthiness and predictive power across applications yang2023context; rezaei2021learning.
Remaining challenges of multimodal AI for scientific discovery lie in integrating heterogeneous data, ranging from atomic-level properties to macroscopic observations, and incorporating synthetic data while addressing the scarcity of high-fidelity datasets amidst an abundance of low-fidelity data. Advancing scientific discovery with multimodal AI requires comprehensive datasets, rigorous data standards, and interdisciplinary research environments. These efforts will ensure that models are aligned with real-world dynamics, unlocking new possibilities in materials science, physics, and beyond trinh2024olympiad; raayoni2021ramanujan.
@@
Multimodal AI can help address sustainability challenges, such as biodiversity conservation and environmental monitoring. Beyond its applications in climate change adaptation, multimodal AI can combine geospatial data, bioacoustic recordings, and environmental DNA to track species and ecosystems pollock2025harnessing, supporting conservation planning and enhancing understanding of ecosystem dynamics. In terrestrial environments, ecoacoustic and LiDAR data can be combined to model biodiversity variation across complex landscapes rappaport2020acoustic, while in marine settings, integrating remote sensing with in situ sensors enables long-term monitoring of ecosystem health and debris pathways maximenko2021integrated.
Geographic imbalances in data availability, particularly in underrepresented regions, limit model fairness and accuracy. Marine environments pose unique sensing challenges briciu2023sensors, including high turbidity and limited optical visibility, which constrain conventional imaging approaches. This necessitates innovative solutions such as autonomous underwater vehicles and acoustic sensors to access deeper waters, remote habitats, and ecologically sensitive regions whitt2020future. Future advancements will depend on building diverse, scalable datasets, de-biasing models, and leveraging digital twins Li2023 for real-time monitoring and actionable insights. By addressing these challenges, multimodal AI can contribute significantly to global sustainability efforts zhao2024artificial.
@@
Multimodal AI can advance finance by integrating financial, environmental, and social data to improve risk assessments, market forecasting, and sustainable investment strategies. In particular, it can enhance ESG (environmental, social, and governance) investing by helping stakeholders make more responsible financial decisions ang2024learning. Integrating financial data and environmental data such as satellite imagery can enable a more comprehensive approach to risk management and resilience leng2024power. Additionally, graph-based multimodal models can monitor systemic risks by capturing complex relationships within financial ecosystems ang2024learning, supporting the early detection of market disruptions. Traditional financial models often struggle to respond to economic shocks such as natural disasters. Integrating real-time, multimodal data can improve model adaptability and accuracy under volatile conditions, ultimately enhancing financial stability cao2023ai.
@@
Challenges such as data privacy, regulatory complexity, and the need for explainability remain significant barriers. Addressing these issues requires comprehensive multimodal datasets and transparent evaluation frameworks to build trust and ensure compliance with regulations. Interdisciplinary collaborations can help refine multimodal AI systems in finance. Environmental science informs environmental risk assessments, enabling more resilient investment strategies and credit evaluations in the face of severe adverse events. Social sciences provide socio-political insights that promote transparency, equity, and ethical compliance, ensuring financial practices align with societal goals. These collaborations align financial systems with ethical and sustainable practices while equipping stakeholders to make resilient decisions in an evolving financial landscape.
Multimodal benchmark datasets
Figure S1: Examples of multimodal benchmark datasets. We selected two illustrative examples for each of the six disciplines to showcase diverse multimodal data and provide a starting point for multimodal AI exploration. For each dataset, we report six key attributes: year of release, data origin, data modalities, accessibility, scale of the primary modalities, and example applications. In each discipline, datasets are listed in ascending order by year of release. For the healthcare, science, and finance disciplines, we selected one small-scale dataset for quick experimentation and one larger-scale dataset for comprehensive exploration. The datasets include: TCGA (The Cancer Genome Atlas) tomczak2015review; MIMIC-IV (The Medical Information Mart for Intensive Care IV database) johnson2023mimic; Trafficking-10k (human trafficking advertisement dataset) tong2017combating; IGDD project (Instagram data donation project) razi2022instagram; nuScenes (autonomous driving scene dataset) caesar2020nuscenes; DAIR-V2X (real-scenarios vehicle to everything dataset) yu2022dair; Materials project (material property dataset) jain2013commentary; MathVista (mathematical reasoning dataset) lu2024mathvista; iNaturalist (citizen science platform for biodiversity data) inaturalist; ERA5 (European Centre for Medium-Range Weather Forecasts Reanalysis v5) hersbach2020era5; Monetary policy calls (monetary policy call dataset) mathur2022monopoly; China A-shares market (dataset of public companies listed in China A-shares market) cheng2022financial. IGDD and ERA5 are not multimodal in the traditional sense but are treated as such under our broader definition.
Benchmark datasets are crucial for developing multimodal AI systems. Figure S1 illustrates examples of selected benchmark datasets across the six discussed disciplines, serving as valuable resources for experimental exploration and evaluation. Each dataset is described by its year of release, data origin, data modalities, accessibility, scale of the primary modalities, and example applications. For the healthcare, science, and finance disciplines, we selected one small-scale dataset for quick experimentation and one larger-scale dataset for comprehensive exploration. In our selection, we prioritise datasets having multiple modalities, particularly those beyond vision and language, and with free and open access, although this is not always possible. For example, the Instagram data donation (IGDD) project and Trafficking-10k datasets require special permissions or approvals. In addition, there are multiple ways to define and categorise the modalities within a dataset, so the modalities listed in Fig. S1 represent just one possible approach. For example, the IGDD dataset contains only text data, and we treat its two subtypes, chat histories and surveys, as distinct text modalities, following our broader definition of multimodality. The ERA5 dataset comprises spatiotemporal data, and we consider variables such as temperature, relative humidity, and wind velocity as multiple ‘views’ (in the broader sense defined in this Perspective), each offering distinct information for weather forecasting and climate research, as in previous studiesjin2023spatiotemporal; chen2023fengwu.
@@
Figure S2: Landscape of multimodal AI research from 2019 to 2024. a, Yearly growth in the number of multimodal AI preprints on arXiv (bars) and their proportion among all AI preprints (line), demonstrating a marked rise from 2019 to 2024. b, Distribution of multimodal AI preprints by the number of combined modalities. The inset line plot shows the growth rate of each combination. c, Detailed breakdown of pairwise, triple, and quadruple modality combinations, demonstrating the dominance of pairwise combinations and the increasing prevalence of more complex combinations. d, Heatmap of combinations of ‘other’ modalities (excluding vision and language) by year. Each row in the heatmap represents the number of preprints published in a given year, with darker shades indicating higher counts. The coloured circles on the left correspond to the modality combinations for each row. For example, in 2023, there were only two preprints using graph and tabular data (the cell in the sixth row and the fifth column).
All arXiv preprints ginsparg2011arxiv from 2019 to 2024 were analysed to examine the latest trends in multimodal AI research. ArXiv was chosen for its open access, broad adoption within the AI community, and ability to reflect emerging developments more rapidly than peer-reviewed platforms. Raw data were extracted from the publicly available arXiv metadata dump arxiv_org_submitters_2024 hosted on Kaggle and filtered using queries for six general AI terms “AI”, “A.I.”, “artificial intelligence”, “machine learning”, “deep learning”, and “neural network” appearing in either the title or the abstract to identify relevant preprints. The list was further refined to identify multimodal AI preprints by searching for “multimodal” and “multi-modal”, and specific modalities were identified through targeted queries: “vision”, “image”, “video”, and “visual” for vision; “text”, “language”, and “textual” for language; and similar sets of terms for the other six modalities: time series, graph, audio, spatial, sensor, and tabular data (see Data availability). As these queries are inherently approximate, conclusions are confined to overall trends supported by strong evidence rather than specific numerical details.
@@
Figures 1 and S2 summarise the findings. Figure 1 shows the dominance of vision and language in multimodal AI research. Figure S2a depicts an accelerating upward trend in both the number and proportion of multimodal AI preprints since 2022, highlighting the field’s rapid expansion. Figure S2b categorises multimodal AI preprints by the number of modalities and their respective growth rates for additional insights. Pairwise combinations are the most popular, while combinations involving more modalities are steadily gaining attention, supporting richer data fusion for tasks requiring diverse information sources. Figure S2c examines pairwise, triple, and quadruple modality combinations across four modality clusters: vision and language, vision and others, language and others, and others. The combination of vision and language remains dominant, particularly in triple and quadruple combinations. In particular, from 2023 to 2024, pairwise combinations involving language have grown significantly, driven by the revolution in large language models. Finally, Fig. S2d details underexplored modality combinations beyond vision and language, ordered by their 2024 preprint counts. Time series and sensor are the most popular non-vision, non-language modality combinations, likely due in part to the fact that sensor data are often recorded as time series. The popularity of the time series and spatial combination may be partially attributed to the importance of spatiotemporal modelling. In contrast, non-vision, non-language modality combinations involving graph, audio, and tabular data are much less explored, highlighting their untapped potential in multimodal AI research.
Figure S3: Deployment-centric multimodal AI workflow for the self-driving car design and climate change adaptation use cases. These examples are depicted in a similar way as Fig. 3, showing the diverse data sources needed and the multidisciplinary collaboration involved.
S3. Illustration of the self-driving car design and climate change adaptation use cases
Figure S3 illustrates the deployment-centric workflow for the self-driving car design and climate change adaptation use cases. In self-driving car design (top), the workflow incorporates deployment constraints from social science, engineering, sustainability, and finance disciplines. Two key tasks focus on engineering and science considerations, respectively. Multimodal AI development uses multiple data sources, including remote sensing, weather, and materials data, aiming for real-world deployment exemplified by automated intersection turning. In climate change adaptation (bottom), the workflow considers constraints such as mass deployment cost barriers, minimum data sampling density, and prioritisation of vulnerable regions. Task formulation and multimodal AI development focus on environmental sustainability and financial considerations, reflecting how task and model development are shaped by discipline-specific needs. Early flood warning serves as an example of real-world deployment and assessment. Together, Fig. 3 and Fig. S3 demonstrate the versatility of the proposed workflow, showing how deployment-centric considerations can be applied across different disciplines to solve complex challenges.

### v2

@@ Acknowledgements
X.L., J.Z., S.Z. and H.L. contributed equally. X.L., J.Z., S.Z., T.L.vdP., S.T., T.Z., W.K.C., P.H.C., and H.L. conceptualised the manuscript. X.L., J.Z., S.Z., A.G., M.Z., L.vZ., O.T., and H.L. designed the figures. X.L. and H.L. coordinated the entire project. All authors contributed to writing, resources, or editing.
@@
Beyond its critical role in pandemic response, multimodal AI holds transformative potential in healthcare by advancing diagnostics, personalising treatment, and improving patient care. By integrating diverse data sources such as electronic health records (EHRs), imaging, biosignals, and multiomics, multimodal AI systems can provide comprehensive insights into patient health Acosta et al. 2022. For example, integrating wearable biosignals with EHRs can enable continuous monitoring, supporting the early detection of chronic conditions such as arrhythmias and heart failure Krittanawong et al. 2021. In oncology, the fusion of imaging modalities and multiomics data can enhance cancer diagnosis and treatment planning Steyaert et al. 2023, offering a more nuanced understanding of disease progression. Similarly, in neurology, combining imaging, sensor data, and clinical observations can facilitate the early detection of neurodegenerative diseases Huang et al. 2023.
Wider adoption of multimodal AI is hindered by challenges such as privacy concerns, the heterogeneity of healthcare data, and the scarcity of high-quality datasets. Addressing these challenges requires standardised methodologies, privacy-preserving techniques, and interdisciplinary collaborations. Building datasets that reflect the complexities of patient pathways and addressing bottlenecks in data integration are crucial steps towards integrating multimodal AI into routine clinical practice. Techniques such as federated learning, differential privacy, and privacy-preserving generative AI for synthetic multimodal data generation can help mitigate privacy risks, enabling the development of robust and trustworthy AI systems. Integrating multimodal data, such as multiomics Lunke et al. 2023, into routine clinical workflows will require rigorous validation and improved standardisation. The limited use of randomised controlled trials (RCTs) to validate AI in clinical settings Han et al. 2024 has prompted the development of community-based guidelines Chan et al. 2020 to improve reliability and transparency.
@@
In social science, multimodal AI can drive innovations in behavioural analysis, public policy decision-making, and societal impact assessments. Beyond its applications in pandemic response, multimodal AI can help monitor societal behaviours and safeguard vulnerable populations, such as protecting children online, by integrating text, image and metadata from social media Ali et al. 2023. Multimodal AI systems can also help detect online criminal activity Goyal et al. 2023 and analyse societal trends to inform public policy Androutsopoulou and Charalabidis 2018. Additionally, acoustic data can add depth to communication analysis by capturing emotional nuances Zhu et al. 2024. Combining economic transaction data with geospatial information can reveal trends in societal decision-making Huang et al. 2024.
@@
Beyond self-driving cars, multimodal AI supports advanced autonomous systems, including robotics, precision manufacturing, and medical applications. Robots equipped with multimodal AI can integrate vision, haptic feedback, and sensor data to perform delicate tasks, such as in surgical procedures He et al. 2022 or complex manipulation Wang et al. 2024, where precision is paramount. Additionally, multimodal AI can enable seamless human-robot interaction by integrating speech, vision, and environmental understanding Duncan et al. 2024. Speech technology and natural language processing can enhance user interfaces, improving interaction while minimising distractions for operators in robotic applications.
Despite the potential of multimodal AI, challenges remain in achieving effective data fusion, system interoperability, and robust model generalisation. Furthermore, the lack of diverse datasets, particularly for edge cases or atypical deployment environments, limits the scalability of these systems. Establishing open platforms for sharing data and toolkits can accelerate innovation and ensure interoperability across engineering applications. Integrating insights from legal studies and social sciences can enhance compliance with evolving regulations and improve user-centred design, boosting societal acceptance of autonomous systems. Environmental science can contribute to the development of energy-efficient and sustainable designs, particularly for robotics and manufacturing processes. Addressing these challenges requires interdisciplinary collaborations to improve data diversity and develop scalable, interoperable systems.
@@
Multimodal AI is transforming scientific discovery by integrating diverse data sources to model complex phenomena. Beyond its role in vehicle dynamics simulation, as seen in self-driving car design, multimodal AI is accelerating advancements in materials science. For instance, integrating experimental and computational data enables the development of improved batteries, fuel cells and supercapacitors Liu et al. 2024. We can incorporate domain knowledge via techniques such as physics-informed neural networks Raissi et al. 2019; Karniadakis et al. 2021 and neurosymbolic AI Garcez and Lamb 2023 to improve the accuracy of scientific simulations and enhance model trustworthiness and predictive power across applications Yang et al. 2023; Rezaei-Shoshtari et al. 2021.
Remaining challenges of multimodal AI for scientific discovery lie in integrating heterogeneous data, ranging from atomic-level properties to macroscopic observations, incorporating synthetic data, and addressing the scarcity of high-fidelity datasets amidst an abundance of low-fidelity data. Advancing scientific discovery with multimodal AI requires comprehensive datasets, rigorous data standards, and interdisciplinary research environments. These efforts will ensure that models are aligned with real-world dynamics, unlocking new possibilities in materials science, physics, and beyond Trinh et al. 2024; Raayoni et al. 2021.
@@
Multimodal AI can help address sustainability challenges, such as biodiversity conservation and environmental monitoring Gui et al. 2024; Van der Plas et al. 2025. Beyond its applications in climate change adaptation, multimodal AI can combine geospatial data, bioacoustic recordings, and environmental DNA to track species and ecosystems Pollock et al. 2025, supporting conservation planning and enhancing understanding of ecosystem dynamics. In terrestrial environments, ecoacoustic and LiDAR data can be combined to model biodiversity variation across complex landscapes Rappaport et al. 2020. In marine settings, integrating remote sensing with in situ sensors enables long-term monitoring of ecosystem health and debris pathways Maximenko et al. 2021.
Geographic imbalances in data availability, particularly in underrepresented regions, limit model fairness and accuracy. Marine environments pose unique sensing challenges Briciu-Burghina et al. 2023, including high turbidity and limited optical visibility, which limit the effectiveness of conventional imaging approaches. This necessitates innovative solutions such as autonomous underwater vehicles and acoustic sensors to access deeper waters, remote habitats, and ecologically sensitive regions Whitt et al. 2020. Future advancements will depend on building diverse, scalable datasets, de-biasing models, and leveraging digital twins Li et al. 2023 for real-time monitoring and actionable insights. By addressing these challenges, multimodal AI can make meaningful contributions to global sustainability efforts Zhao et al. 2024.
@@
Multimodal AI can advance finance by integrating financial, environmental, and social data to improve risk assessments, market forecasting, and sustainable investment strategies. In particular, it can enhance ESG (environmental, social, and governance) investing by helping stakeholders make more responsible financial decisions Ang and Lim 2024. Integrating financial data and environmental data, such as satellite imagery, can enable a more comprehensive approach to risk management and resilience Leng et al. 2024. Additionally, graph-based multimodal models can monitor systemic risks by capturing complex relationships within financial ecosystems Ang and Lim 2024, supporting the early detection of market disruptions. Traditional financial models often struggle to respond to economic shocks such as natural disasters. Integrating real-time, multimodal data can improve model adaptability and accuracy under volatile conditions, ultimately enhancing financial stability Cao 2023.
@@
Challenges such as data privacy, regulatory complexity, and the need for explainability continue to be substantial barriers. Addressing these issues requires comprehensive multimodal datasets and transparent evaluation frameworks to build trust and ensure compliance with regulations. Interdisciplinary collaborations can help refine multimodal AI systems in the finance sector. Environmental science informs environmental risk assessments, enabling more resilient investment strategies and credit evaluations in the face of severe adverse events. Social sciences provide socio-political insights that promote transparency, equity, and ethical compliance, ensuring that financial practices align with societal goals. These collaborations align financial systems with ethical and sustainable practices while equipping stakeholders to make resilient decisions in an evolving financial landscape.
@@
Supplementary Figure 1: Landscape of multimodal AI research (2019–2024). a, Yearly growth in the number of multimodal AI preprints on arXiv (bars) and their proportion among all AI preprints (line), both showing an accelerating upward trend since 2022, reflecting the field’s rapid expansion, likely driven by the large language model (LLM) revolution. b, Distribution of multimodal AI preprints by the number of combined modalities. The inset line plot shows the growth rate of each combination. While pairwise combinations remain the most prevalent, combinations involving more modalities are steadily gaining attention, reflecting increasing interest in richer data fusion for tasks requiring diverse information sources. c, Detailed breakdown of pairwise, triple, and quadruple modality combinations, highlighting trends across four modality clusters: vision and language, vision and other(s), language and other(s), and others. Combinations involving vision and language remain dominant. In particular, the number of pairwise combinations involving language has increased substantially from 2023 to 2024, likely due to the LLM-driven surge in multimodal research.
All arXiv preprints Ginsparg 2011 from 2019 to 2024 were analysed to examine the latest trends in multimodal AI research. ArXiv was chosen for its open access, broad adoption within the AI community, and ability to reflect emerging developments more rapidly than peer-reviewed platforms. Raw data were extracted from the publicly available arXiv metadata dump arXiv.org submitters 2024 hosted on Kaggle and filtered using queries for six general AI terms “AI”, “A.I.”, “artificial intelligence”, “machine learning”, “deep learning”, and “neural network” appearing in either the title or the abstract to identify relevant preprints. The list was further refined to identify multimodal AI preprints by searching for “multimodal” and “multi-modal”, and specific modalities were identified through targeted queries: “vision”, “image”, “video”, and “visual” for vision; “text”, “language”, and “textual” for language; and similar sets of terms for the other six modalities: time series, graph, audio, spatial, sensor, and tabular data (see Data availability). As these queries are inherently approximate, conclusions are confined to overall trends supported by strong evidence rather than specific numerical details.
Supplementary Figure 1 provides more detailed analyses of the multimodal AI landscape, including statistics on triple and quadruple modality combinations.